# A Spatial Gap in the Sky Distribution of Fast Radio Burst Detections Coinciding with Galactic Plasma Overdensities

### v1

@@ 1 Introduction
Fast radio bursts (FRBs) are bright and brief extragalactic bursts of energy emitted in the radio spectrum (D. R. Lorimer et al., 2007, also see E. Petroff et al., 2022 for a recent overview). While their origins remain unclear, there are a growing number of observations that support models based on compact objects, such as highly magnetized neutron stars (e.g.CHIME/FRB Collaboration et al., 2020; C. D. Bochenek et al., 2020, see E. Platts et al., 2019 for a collection of proposed models). Regardless of their origins, FRBs undergo the same dispersive and scatter-broadening effects regularly observed in Galactic radio pulsars (S. K. Ocker et al., 2022), with scattering arising from dense ionized regions in near-source environments within their host galaxies and/or and the Milky Way. Scattering often needs to be modeled with one or more distinct scattering screens – a framework developed for pulsars (I. P. Williamson, 1972) and subsequently applied to FRBs. In this sense, FRBs can be used as powerful probes of the electron density fluctuations in the interstellar medium (ISM) along their lines of sight.
@@
Along a given line of sight (LOS), the scattering timescale ($\tau_{\rm sc}$) characterizes the time delay caused by multi-path propagation through the warm ionized medium (WIM). Propagation through ionized plasma also results in a frequency-dependent dispersive delay that is proportional to the integrated column density of free electrons along the LOS. This integrated quantity is known as the dispersion measure (DM). Based on the DM and $\tau_{\rm sc}$ predictions from the NE2001 Galactic electron density model (J. M. Cordes & T. J. W. Lazio, 2002), S. K. Ocker et al. (2022) postulated a “zone of avoidance" to span over a range of Galactic longitudes $|l|\leq 50\degree$ and latitudes $|b|\leq 4.1\degree$ at 0.4 GHz for FRBs. J.-P. Macquart & S. Johnston (2015) proposed that diffractive scintillation enhances FRB detectability at higher Galactic latitudes by boosting some bursts above the detection threshold, particularly in the presence of a steep luminosity distribution. Strong scattering in the Galactic plane can suppress scintillation and broaden the pulse, reducing the likelihood of such boosts.
@@
Studies of the FRB sky distribution have been limited by the small number of detections and the challenges associated with combining data from different instruments, each with differing sensitivities, instrumental uncertainties, and selection functions. The FRB sky distribution can greatly vary based on the observing frequency. As a result, the scope of such studies so far has been limited to probing only the Galactic latitude dependence. Using data acquired with the Canadian Hydrogen Intensity Mapping Experiment (CHIME) Fast Radio Burst project (hereafter, CHIME/FRB), A. Josephy et al. (2021) reported no significant dependence of FRB detections on Galactic latitude in CHIME/FRB Catalog 1 (CHIME/FRB Collaboration et al. 2021) containing 536 FRBs.
@@
The forthcoming second catalog of CHIME/FRB (hereafter Catalog 2) represents a significant advancement, reporting the largest number of unique FRBs detected by a single instrument (CHIME/FRB Collaboration et al. 2025). An updated map of the FRB sky positions in Catalog 2 reveals a striking and visually discernible gap in FRB detections. For the first time, we clearly identify a Galactic scattering zone of avoidance (hereafter referred to as the “detection gap”) in FRB detections, which spans Galactic longitudes $70\degree<l<90\degree$ and latitudes $7\degree<b<11\degree$ in the CHIME observing band (400–800 MHz). This detection gap coincides with the Cygnus X region – a massive star-forming region in the Galaxy, known to contain dense ionized gas and strong degrees of turbulence (J. H. Piddington & H. C. Minnett, 1952a; H. J. Wendker et al., 1991; N. Schneider et al., 2006). The alignment with known Galactic structure suggests a strong link between plasma propagation effects and the suppression of FRB detectability in the CHIME/FRB experiment.
@@
This study reports a detailed analysis of the detection gap, investigates the extent to which propagation effects contribute to it and identifies associated patterns in scattering and detectability. The structure of the paper is as follows. In §2, we describe the observational dataset including post-detection selection criteria and classification of FRBs based on morphology. In §3 we present evidence for a significant detection gap observed in the sky distribution of CHIME/FRB Catalog 2. In §4, we explore potential observational and astrophysical explanations for the detection gap, and derive an empirical lower bound on $\tau_{\rm sc}$ responsible for it. The role of Galactic scattering in shaping FRB detectability across the full Catalog 2 sample is discussed in §5. Our conclusions are summarized in §6.
@@
2 Observations
In this work, we analyze data products generated by the CHIME/FRB backend for Catalog 2 (CHIME/FRB Collaboration et al. 2025). Full descriptions of the CHIME telescope and its radio-transient backends are provided in other works ($al@abb+18,abb+21,abb+22; \al@abb$+18,abb+21,abb+22). We nonetheless summarize aspects of CHIME and the Catalog 2 data products relevant to our study in this section.
2.1 The CHIME Telescope and Instrumentation
@@
2.2 FRB Celestial Positions
[…]
2.3 Model-dependent Classification of FRBs
@@
Statistically significant signals detected by the CHIME/FRB backend are classified as “Galactic”, “ambiguous”, or “extragalactic”, based on the source position and DM estimated by the real-time bonsai tree de-dispersion algorithm (CHIME/FRB Collaboration et al. 2018). Single pulses from Galactic radio pulsars are identified through coincidences in both position and DM, and for the purposes of Catalog 2 are excluded from subsequent analysis. Our signals of interest are those deemed extragalactic (i.e., FRBs), which are classified as such if their DMs exceed both of the predicted maximum values from the NE2001 (J. M. Cordes & T. J. W. Lazio, 2002) and YMW16 (J. M. Yao et al., 2017) electron density models for their lines of sight; these predictions are made assuming a distance of 25 kpc between the observer and the edge of the Milky Way disk. No aspects of scatter-broadening are considered for the real-time classification of FRBs.
2.4 Best-fit Models of FRB Morphology
@@
In order to characterize the effects of the WIM on burst morphology in Catalog 2, we use spectro-temporal measurements of FRB morphology estimated with fitburst (E. Fonseca et al., 2024). The fitburst modeling framework assumes that each pulsed feature undergoes cold-plasma dispersion and, in the absence of scattering-broadening, has a Gaussian shape of intrinsic temporal width ($\sigma$). If relevant, the shape of the burst is instead assumed to be the pulse broadening function of a Gaussian profile (e.g., M. M. McKinnon, 2014) that depends on $\sigma$ and $\tau_{\rm sc}$.
@@
All Catalog 2 fitburst measurements assume that $\tau_{\rm sc}$ scales with electromagnetic frequency ($\nu$) such that $\tau_{\rm sc}\propto\nu^{-4}$. The Catalog 2 values for $\tau_{\rm sc}$ are originally reported referenced to $\nu=400.195$ MHz (CHIME/FRB Collaboration et al. 2025); however, for this present study we scale $\tau_{\rm sc}$ to be referenced at $\nu=1$ GHz using $\tau_{\rm sc,\,\nu}\propto\nu^{-4}$ (N. D. R. Bhat et al., 2004), as is typically done for related studies on radio pulsars and FRBs. Additionally, for detectability discussions, we reference $\tau_{\rm sc}$ to CHIME frequencies whenever relevant – explicitly denoting it as $\tau_{\rm sc,\,600\,MHz}$. We used $\tau_{\rm sc,\,1\,GHz}>0.13$ ms as a threshold for determining which FRBs have “well-measured" scatter-broadening; this threshold value was obtained through comparisons of fitburst fits between total-intensity and microsecond-resolution baseband (voltage) data of the same detections (K. R. Sand et al., 2025).
2.5 Post-detection Selection Criteria for Catalog 2 FRBs
@@
We adopted the initial data quality flags outlined in the Catalog 2 methodology (CHIME/FRB Collaboration et al. 2025). In addition, we required that the event had successful fits from the header_localization and fitburst pipelines. We excluded FRBs detected in the sidelobes of the primary beam, since these are more complicated to localize accurately (H.-H. Lin et al., 2024). The above selection criteria resulted in 4439 FRBs. To obtain the full set of unique sightlines from Catalog 2, we retain only the burst with the highest $(S/N)$ for each repeating source. The final sample utilized in this study consists of 3552 FRBs – comprising 3469 non-repeaters and 83 unique sightlines from repeating FRBs. Of those, 1146 FRBs meet the criterion for well-measured $\tau_{\rm sc,\,1\,GHz}$ described in §2.4.
3 A Detection Gap in the sky distribution of FRBs
@@
Figure 1 presents the spatial distribution of FRBs in Catalog 2, with the detection gap region highlighted by a cyan dashed enclosure. The high density of FRB detections near the North Celestial Pole is due to CHIME’s declination-dependent sensitivity and exposure. A magnified view of the detection gap region is presented in Figure 2. To delineate its shape and extent, we applied geometric methods based on the Delaunay triangulation and its dual, the Voronoi diagram using the scipy.spatial module (P. Virtanen et al., 2020). Using the Voronoi diagram, we identified the largest empty circle within the detection gap by locating the Voronoi vertices farthest from any FRB. We estimate the center of the largest empty circle (LEC) within the detection gap to be at $(l,b)=(77.7\degree,0.9\degree)$, with a radius of $R_{\rm LEC}=7.7\degree$. The triangulation partitions FRB positions into non-overlapping triangles such that no point is inside the circumcircle of any triangle. We identified the triangles for which the distance between their circumcenters and the center of gap fell within the radius. The convex hull obtained from connecting the outer boundary of these triangles is the polygon that encloses the sky with zero FRB detections in Figure 2, with an area of 216.2 deg2.
@@
At 600 MHz, the effective gain of CHIME along the detection gap LOS is $G\approx 1.02$ K/Jy with system temperature $T_{\rm sys}\approx$ 50 K (CHIME/FRB Collaboration et al., 2018). Thus, the system equivalent flux density (SEFD) is roughly $T_{\rm sys}/G\approx$ 48.78 Jy. The single-pulse radiometer equation gives a minimum detectable flux density of $\sim 0.5$ Jy for narrow pulses ($\sim 1$ ms broadened width) and $\sim 0.1$ Jy for pulses with broadened width larger than 10 ms along this LOS. In the following discussion, we assess whether the apparent detection gap can be explained by instrumental sensitivity and statistical noise effects, or whether it points to suppression in detectability due to scattering caused by ionized Galactic structures.
3.1 Statistical Significance of the Gap
@@
To evaluate the significance of the lack of detections, we compute the probability of detecting 0 FRBs within a circle of radius $R_{\rm LEC}$ or larger anywhere on the sky surveyed by CHIME/FRB assuming a non-homogeneous Poisson process model for its detections. The non-homogeneous Poisson process is described by a position-dependent intensity function $\lambda(\alpha,\delta)$ which here represents the expected density of detected FRBs and scales with the instrument’s exposure and sensitivity. It is convenient to work in the equatorial coordinate system ($\alpha,\delta$) as the CHIME/FRB gain varies primarily with $\delta$. Moreover, the cosine dependence of an isotropic universal distribution of FRBs on the celestial sphere cancels out with the inverse-cosine dependence of the exposure in declination.
@@
As reported by CHIME/FRB Collaboration et al. (2021), we note reduced exposure between $27\degree<\delta<34\degree$ as shown in Figure 8 in Appendix C and Figure 5 of CHIME/FRB Collaboration et al. 2021. This is due to a time-limited failure of one of the four CPU nodes covering these declinations. The patch of reduced exposure overlaps with only 7% of the total sky area of the detection gap. Omitting the reduced exposure patch, we fit a polynomial to the logarithm of the exposure from Catalog 2 versus $\delta$ and find that the observed exposure in this region is $\approx 18$% lower than expected. At $\delta>70\degree$, sources are circumpolar from CHIME’s latitude and cross the telescope’s meridian twice daily. These ’upper’ and ’lower’ transits therefore yield higher effective exposure at $\delta>70\degree$. Using the CHIME/FRB sensitivity approximation parameterized by A. M. Cook et al. (2024) and accounting for exposure as described above, we can write
@@
where $H(x)$ is the Heaviside step function, $\gamma=1.5$ is the Euclidean power-law index for a flux-limited survey representing the FRB luminosity function, $\phi=49.32\degree$ is the geographic latitude of the CHIME telescope (CHIME/FRB Collaboration et al. 2018), and $A$ is the following normalization constant such that the intensity equals $N_{\rm total}$, the total number of FRBs in the sample, when integrated over CHIME/FRB’s sky:

### v2

@@ I Introduction
Fast radio bursts (FRBs) are bright and brief extragalactic bursts of energy emitted in the radio spectrum (lbm+07, also see petroff_frb_overview for a recent overview). While their origins remain unclear, there are a growing number of observations that support models based on compact objects, such as highly magnetized neutron stars (e.g.chime_frb_galactic_magnetar; brb+20, see pww+19 for a collection of proposed models). Regardless of their origins, FRBs undergo the same dispersive and scatter-broadening effects regularly observed in Galactic radio pulsars (occ_2022), with scattering arising from dense ionized regions in near-source environments within their host galaxies and/or and the Milky Way. Scattering often needs to be modeled with one or more distinct scattering screens – a framework developed for pulsars (Williamson_mnras) and subsequently applied to FRBs. In this sense, FRBs can be used as powerful probes of the electron density fluctuations in the interstellar medium (ISM) along their lines of sight.
@@
Along a given line of sight (LOS), the scattering timescale ($\tau_{\rm sc}$) characterizes the time delay caused by multi-path propagation through the warm ionized medium (WIM) of the Milky Way Galaxy. Propagation through ionized plasma also results in a frequency-dependent dispersive delay that is proportional to the integrated column density of free electrons along the LOS. This integrated quantity is known as the dispersion measure (DM). Based on the DM and $\tau_{\rm sc}$ predictions from the NE2001 Galactic electron density model (ne2001), occ_2022 postulated a “zone of avoidance" to span over a range of Galactic longitudes $|l|\leq 50\degree$ and latitudes $|b|\leq 4.1\degree$ at 0.4 GHz for FRBs. mj15 proposed that diffractive scintillation enhances FRB detectability at higher Galactic latitudes by boosting some bursts above the detection threshold, particularly in the presence of a steep luminosity distribution. Strong scattering in the Galactic plane can suppress scintillation and broaden the pulse, reducing the likelihood of such boosts.
@@
FRBs are therefore understood to encode information about the WIM through measurements of $\tau_{\rm sc}$, serving as “backlights" to the foreground that is the Milky Way Galaxy. A large sample of $\tau_{\rm sc}$ measured from FRBs observed across a large fraction of the sky are expected to provide the means for constraining structural information of the WIM (e.g., mj15). However, past studies of the FRB sky distribution have been limited by the small number of detections and the challenges associated with combining data from different instruments, each with differing sensitivities, instrumental uncertainties, and selection functions. The FRB sky distribution can greatly vary based on the observing frequency. As a result, the scope of such studies so far has been limited to probing only the Galactic latitude dependence (phl19). Using data acquired with the Canadian Hydrogen Intensity Mapping Experiment (CHIME) Fast Radio Burst project (hereafter, CHIME/FRB), jcc+21 reported no significant dependence of FRB detections on Galactic latitude in CHIME/FRB Catalog 1 (chimefrbcatalog1) containing 536 FRBs.
@@
The forthcoming second catalog of CHIME/FRB (hereafter Catalog 2) represents a significant advancement, reporting the largest number of unique FRBs detected by a single instrument (chimefrbcatalog2). An updated map of the FRB sky positions in Catalog 2 reveals a striking and visually discernible gap in FRB detections. For the first time, we clearly identify a Galactic scattering zone of avoidance (hereafter referred to as the “detection gap”) in FRB detections, which spans Galactic longitudes $70\degree<l<90\degree$ and latitudes $7\degree<b<11\degree$ in the CHIME observing band (400–800 MHz). This detection gap coincides with the Cygnus X region – a massive star-forming region in the Galaxy, known to contain dense ionized gas and strong degrees of turbulence (pm52; wendker; schneider+06). The alignment with known Galactic structure suggests a strong link between plasma propagation effects and the suppression of FRB detectability in the CHIME/FRB experiment.
@@
This study reports a detailed analysis of the detection gap, investigates the extent to which propagation effects contribute to it and identifies associated patterns in scattering and detectability. The structure of the paper is as follows. In §II, we describe the observational dataset including post-detection selection criteria and classification of FRBs based on morphology. In §III we present evidence for a significant detection gap observed in the sky distribution of CHIME/FRB Catalog 2. In §IV, we explore potential observational and astrophysical explanations for the detection gap, and derive an empirical lower bound on $\tau_{\rm sc}$ responsible for it. The role of Galactic scattering in shaping FRB detectability across the full Catalog 2 sample is discussed in §V. Our conclusions are summarized in §VI.
Figure 1: Positions of 3542 FRBs in our sample overlaid with the Planck emission measure map (planckforegrounds16). We qualitatively see an anti-correlation between FRB detections and regions of high EM. The detection gap is highlighted with dashed cyan lines. A zoomed-in view of this region is presented in Figure 3.
II Observations
In this work, we analyze data products generated by the CHIME/FRB backend for Catalog 2 (chimefrbcatalog2). Full descriptions of the CHIME telescope and its radio-transient backends are provided in other works (abb+18; abb+21; abb+22). We nonetheless summarize aspects of CHIME and the Catalog 2 data products relevant to our study in this section.
II.1 The CHIME Telescope and Instrumentation
@@
II.2 FRB Celestial Positions
@@
II.3 Model-dependent Classification of FRBs
@@
Statistically significant signals detected by the CHIME/FRB backend are classified as “Galactic”, “ambiguous”, or “extragalactic”, based on the source position and DM estimated by the real-time bonsai tree de-dispersion algorithm (abb+18). Single pulses from Galactic radio pulsars are identified through coincidences in both position and DM, and for the purposes of Catalog 2 are excluded from subsequent analysis. Our signals of interest are those deemed extragalactic (i.e., FRBs), which are classified as such if their DMs exceed both of the predicted maximum values from the NE2001 (ne2001) and YMW16 (ymw17) electron density models for their lines of sight; these predictions are made assuming a distance of 25 kpc between the observer and the edge of the Milky Way disk. No aspects of scatter-broadening are considered for the real-time classification of FRBs.
II.4 Best-fit Models of FRB Morphology
@@
In order to characterize the effects of the WIM on burst morphology in Catalog 2, we use spectro-temporal measurements of FRB morphology estimated with fitburst (fpb+24). The fitburst modeling framework assumes that each pulsed feature undergoes cold-plasma dispersion and, in the absence of scattering-broadening, has a Gaussian shape of intrinsic temporal width ($\sigma$). If relevant, the shape of the burst is instead assumed to be the pulse broadening function of a Gaussian profile (e.g., mck14) that depends on $\sigma$ and $\tau_{\rm sc}$.
@@
All Catalog 2 fitburst measurements assume that $\tau_{\rm sc}$ scales with electromagnetic frequency ($\nu$) such that $\tau_{\rm sc}\propto\nu^{-4}$. The Catalog 2 values for $\tau_{\rm sc}$ are originally reported referenced to $\nu=400.195$ MHz (chimefrbcatalog2); however, for this present study we scale $\tau_{\rm sc}$ to be referenced at $\nu=1$ GHz using $\tau_{\rm sc,\,\nu}\propto\nu^{-4}$ (bcc+04), as is typically done for related studies on radio pulsars and FRBs. Additionally, for detectability discussions, we reference $\tau_{\rm sc}$ to CHIME frequencies whenever relevant – explicitly denoting it as $\tau_{\rm sc,\,600\,MHz}$. We used $\tau_{\rm sc,\,1\,GHz}>0.13$ ms as a threshold for determining which FRBs have “well-measured" scatter-broadening; this threshold value was obtained through comparisons of fitburst fits between total-intensity and microsecond-resolution baseband (voltage) data of the same detections (sand+25_published).
II.5 Post-detection Selection Criteria for Catalog 2 FRBs
@@
We adopted the initial data quality flags outlined in the Catalog 2 methodology (chimefrbcatalog2). In addition, we required that the event had successful fits from the header_localization and fitburst pipelines. We excluded FRBs detected in the sidelobes of the primary beam, since these are more complicated to localize accurately (lsn+23). The above selection criteria resulted in 4429 FRBs. To obtain the full set of unique sightlines from Catalog 2, we retain only the burst with the highest $(S/N)$ for each repeating source. The final sample utilized in this study consists of 3542 FRBs – comprising 3459 non-repeaters and 83 unique sightlines from repeating FRBs. Of those, 1105 FRBs meet the criterion for well-measured $\tau_{\rm sc,\,1\,GHz}$ described in §II.4.
III Statistical Significance of the Detection Gap
@@
Figure 1 presents the spatial distribution of FRBs in Catalog 2, with the detection gap region highlighted by a cyan dashed enclosure. The high density of FRB detections near the North Celestial Pole is due to CHIME’s declination-dependent sensitivity and exposure. A magnified view of the detection gap region is presented in Figure 3. To delineate its shape and extent, we applied geometric methods based on the Delaunay triangulation and its dual, the Voronoi diagram using the scipy.spatial module (scipy). Using the Voronoi diagram, we identified the largest empty circle within the detection gap by locating the Voronoi vertices farthest from any FRB. We estimate the center of the largest empty circle (LEC) within the detection gap to be at $(l,b)=(77.7\degree,0.9\degree)$, with a radius of $R_{\rm LEC}=7.7\degree$. The triangulation partitions FRB positions into non-overlapping triangles such that no point is inside the circumcircle of any triangle. We identified the triangles for which the distance between their circumcenters and the center of gap fell within the radius. The convex hull obtained from connecting the outer boundary of these triangles is the polygon that encloses the sky with zero FRB detections in Figure 3, with an area of 213.6 deg2.
@@
At 600 MHz, the effective gain of CHIME along the detection gap LOS is $G\approx 1.02$ K/Jy with system temperature $T_{\rm sys}\approx$ 50 K (abb+18). Thus, the system equivalent flux density (SEFD) is roughly $T_{\rm sys}/G\approx$ 48.78 Jy. The single-pulse radiometer equation gives a minimum detectable flux density of $\sim 0.5$ Jy for narrow pulses ($\sim 1$ ms broadened width) and $\sim 0.1$ Jy for pulses with broadened width larger than 10 ms along this LOS. In the following discussion, we assess whether the apparent detection gap can be explained by instrumental sensitivity and statistical noise effects, or whether it points to suppression in detectability due to scattering caused by ionized Galactic structures.
Figure 2: Distribution of $R_{\rm LEC}$ estimates obtained from a Delauney-triangulation analysis of the Catalog 2 data (red) and simulations of sky positions of FRBs in $N_{\rm sim}=10^{4}$ catalogs derived from the statistics of Catalog 2 detections (blue-filled). The black-dashed vertical line indicates the largest measured $R_{\rm LEC}$ in Catalog 2.
We use two methods to evaluate the significance of the lack of detections. For the first method, we estimate the probability of obtaining detection gaps with sizes $R_{\rm LEC}\geq 7.7^{\circ}$ assuming that no “foreground" process produces such gaps. We determine this probability by first using Catalog 2 astrometric data to derive a two-dimensional probability density distribution of FRB detections across the CHIME sky through Gaussian kernel-density estimation. We then resample this distribution to simulate a catalog of a size equal to that of Catalog 2. We then execute the Delauney-triangulation algorithm described above on the subset of simulated FRBs within the range $30^{\circ}<b<30$ and $50^{\circ}<l<200^{\circ}$ to find gaps in this realization. We repeat this process $N_{\rm sim}=10^{4}$ times in order to estimate a “realization-averaged" distribution of $R_{\rm LEC}$, which is shown in Figure 2. Using this approach, we find that the probability of obtaining detection gaps with sizes $R_{\rm LEC}\geq 7.7^{\circ}$ due to statistical fluctuations in our measurements is $p\approx 1.3\times 10^{-6}.$
@@
For the second method, we compute the probability of detecting 0 FRBs within a circle of radius $R_{\rm LEC}$ or larger anywhere on the sky surveyed by CHIME/FRB assuming a non-homogeneous Poisson process model for its detections. The non-homogeneous Poisson process is described by a position-dependent intensity function $\lambda(\alpha,\delta)$ which here represents the expected density of detected FRBs and scales with the instrument’s exposure and sensitivity. It is convenient to work in the equatorial coordinate system ($\alpha,\delta$) as the CHIME/FRB gain varies primarily with $\delta$. Moreover, the cosine dependence of an isotropic universal distribution of FRBs on the celestial sphere cancels out with the inverse-cosine dependence of the exposure in declination.
@@
As reported by chimefrbcatalog1, we note reduced exposure between $27\degree<\delta<34\degree$ as shown in Figure 9 in Appendix C and Figure 5 of chimefrbcatalog1. This is due to a time-limited failure of one of the four CPU nodes covering these declinations. The patch of reduced exposure overlaps with only 7% of the total sky area of the detection gap. Omitting the reduced exposure patch, we fit a polynomial to the logarithm of the exposure from Catalog 2 versus $\delta$ and find that the observed exposure in this region is $\approx 18$% lower than expected. At $\delta>70\degree$, sources are circumpolar from CHIME’s latitude and cross the telescope’s meridian twice daily. These ‘upper’ and ‘lower’ transits therefore yield higher effective exposure at $\delta>70\degree$. Using the CHIME/FRB sensitivity approximation parameterized by 2024arXiv241012146C and accounting for exposure as described above, we can write
@@
where $H(x)$ is the Heaviside step function, $\gamma=1.5$ is the Euclidean power-law index for a flux-limited survey representing the FRB luminosity function, $\phi=49.32\degree$ is the geographic latitude of the CHIME telescope (abb+18), and $A$ is the following normalization constant such that the intensity equals $N_{\rm total}$, the total number of FRBs in the sample, when integrated over CHIME/FRB’s sky: